# Adaptive Experiments Toward Learning Treatment Effect Heterogeneity

### v2

@@ 1.1 Motivation
Existing literature in this research area mostly focuses on conducting retrospective analyses employing observational or randomized experiment data. Even with large-scale observational data or carefully collected randomized experiment data, statistical bias in these analyses cannot be overlooked. On the one hand, in observational studies, statistical bias can arise due to potential violations of untestable causal assumptions [4, 5, 11, 16, 31, 39, 50, 60]. For example, one of the commonly imposed causal assumptions in practice is the unconfoundedness assumption, which states that conditional on measured confounders, the treatment assignment is as good as random [see 26, for example]. Due to the untestable nature of the unconfoundedness assumption and the possibility of unmeasured confounders in observational data, the validity of established causal conclusions under this assumption cannot be guaranteed [40, 50]. On the other hand, when analyzing classical randomized experiment data, although carrying out valid causal conclusion does not require imposing untestable causal assumptions, exploring treatment effect heterogeneity could still be susceptible to the winner’s curse bias if seemingly promising heterogeneous treatment effects are selected from the data in an ad hoc fashion [2, 28, 51, 75].
@@ 1.2 Our contribution
We tackle two major challenges in our theoretical investigations, and the techniques developed can be potentially applied in more general design frameworks. First, as the proposed allocation rule depends on the (unknown) mean and variance of the potential outcomes, a critical step is to show that these unknown features are consistently estimable. This is considerably more difficult compared to the i.i.d. setting: our data is adaptively accumulated according to a sequentially estimated and revised treatment assignment rule; hence, observations are, in general, correlated. Leveraging upon martingale methods, we develop a general framework and show that consistent estimation of features of the potential outcome distributions is possible for a large class of adaptive experimental designs (see Lemma C.1 and its corollary in the Supplementary Materials). To be more precise, this result not only unifies adaptive enrichment designs and response adaptive randomization but also applies to other adaptive design settings with objective functions different from ours. We expect the general consistency result to be of independent interest. The second challenge in our theoretical investigation is that the optimized treatment assignment rule, both in finite samples and asymptotically, may not be unique, rendering standard M-estimation proof techniques inapplicable to our setting. Nevertheless, we are able to show that with probability approaching one, any empirically optimized treatment assignment rule will be arbitrarily close to one of the oracle allocations (Lemma C.2 in the Supplementary Materials). Such “set consistency” result seems new in the literature on adaptive experimental design.
@@ 1.3 Existing literature
Our framework falls into the realm of frequentist response adaptive designs. Such design strategies can be roughly divided into response-adaptive randomization (RAR) design and adaptive enrichment design in the existing literature. RAR design often refers to the design strategy in which the treatment assignment probabilities are adapted during the experiment based on the accrued evidence in the outcomes, with the goal of simultaneously achieving the experimental objectives and preserving statistical inference validity [35, 64, 62]. The classical RAR framework often revises the treatment assignment probabilities at infinitely many stages, a design strategy we refer to as fully adaptive settings. In fully adaptive settings, a popular class of response-adaptive randomization designs is the doubly adaptive biased coin (DBCD) design. The early DBCD design can be found in [21], which has its root in Efron’s biased coin design [19]. The asymptotic properties of DBCD designs are studied in various works [36, 37, 78, 96]. Other than the fully adaptive settings, existing RAR designs also accommodate multi-stage settings [58, 92]. Such multi-stage designs propose to revise the treatment assignment probability by minimizing the asymptotic variance of the average treatment effect estimator. Nevertheless, this design is carried out in two stages and is not designed to identify treatment effect heterogeneity.
Instead of relying solely on the outcome variable to optimize for the experimental goals in RAR, one may further incorporate covariate information. RAR that further incorporate covariate information is known as covariate-adjusted response-adaptive (CARA) designs [7, 67, 91]. Early work in [59] and [89] proposes to balance covariates based on the biased coin design. [38] propose a family of CARA designs that could account for both efficiency and ethics. [95] generalizes CARA to incorporate semiparametric estimates. Some related CARA designs are also discussed in [49, 79, 93, 92]. We note that much of the existing work on response-adaptive randomization designs and covariate-adjusted response-adaptive aims to optimize the estimation efficiency of the overall treatment effect but is not tailored to study treatment effect heterogeneity [34, 65].
@@ 4.1 Oracle treatment allocation in response-adaptive randomization designs
This equation suggests that the required number of participants in the treatment arm of subgroup $j$ is reduced when it is relatively easier to distinguish subgroup $j$ from the best subgroup. This occurs when there is a larger difference between $\tau_{j}$ and $\tau_{1}$, or when the variance of subgroup $j$ is higher. To provide a clearer understanding, we consider a simple scenario with $m=3$. In Figure 1, we plot the relationship between $e^{*}_{2}$, $\tau_{2}$, and $\sigma_{2}(1)^{2}$.
@@ 5.1 Oracle subgroup enrichment proportions
In enrichment designs, the treatment allocation probabilities are often fixed, we thus denote the rate function as a function of subgroup enrollment proportions:
@@ 5.2 Proposed adaptive enrichment designs
Applying the same reasoning as discussed in Section 4.3, we calculate the estimated optimal subgroup enrichment proportion by subtracting the proportion of enrolled subjects in the previous stages. If Eq (11) suggests multiple possible designs, one can either choose an enrichment proportion that is closest to the uniform allocation $1/m$ or choose an enrichment proportion that is higher for the subgroup with a larger treatment effect.
@@
Construct the final subgroup treatment effect estimator $\widehat{\tau}_{j}^{AE}$ and the variance estimator $\widehat{\mathbb{V}}_{j}^{AE}$ under the adaptive enrichment (“AE”) design. These two estimators take the same form as in Eq (5) and Eq (6). Then, identify the best subgroup as the one exhibiting the maximal treatment effect size:
@@
$\Bigg{[}\widehat{\tau}^{\texttt{AE}}_{j^{*}}\pm\Phi^{-1}(1-\alpha/2)\cdot\sqrt{\widehat{\mathbb{V}}^{\texttt{AE}}_{j^{*}}/N}\Bigg{]},\ j=1\ldots,m.$
@@ 6.3 Comparison with completely randomized experiments
[…]
@@ 7.2 Synthetic data generation and simulation setup
Under each design setup, we compare our proposed design strategy with other conventional designs as summarized in Table 2. In Table 2 , the “complete randomization” design refers to setting the treatment assignment probability $e_{tj}=\frac{1}{2}$ in all experimental stages, for $t=1,\ldots,T,\ j=1,\ldots,m$. The “doubly adaptive biased coin design” [90] is a conventional response-adaptive randomization design. The “equal enrichment design” refers to the design that sets the enrichment proportion as $p_{tj}=\frac{1}{m}$ across all the experimental stages. The “adaptive enrichment with combination testing” approach is a method that distributes the type I error rate across experimental stages. Based on the computed type I error rate each stage aims to reach, the corresponding enrichment proportions can be estimated. We implement the combination testing approach using R package rpact [46].
@@
(c) Doubly adaptive biased coin design
@@ 7.3 Synthetic case study results
In Figures 4 to 6, we compare our proposed response-adaptive randomization design with the other conventional designs in the fully adaptive setting and the multi-stage setting. We summarize our simulation results from three aspects, following the order outlined at the start of Section 7
@@
Figure 4: The estimated treatment effects and the associated standard errors in the four clothing categories under the complete randomization design, the doubly adaptive biased coin design, and our proposed response-adaptive randomization design in the fully adaptive setting ($N=400$).
Figure 5: Comparison of the proposed response-adaptive randomization design, the complete randomization design, and the doubly adaptive biased coin design under the fully adaptive setting. (A) shows the sample size comparison under various correct selection probability levels. (B) shows the $\sqrt{N}$-scaled winner’s curse bias comparison with respect to different sample sizes.
@@ 8 Discussion
In this manuscript, we propose a unified adaptive experimental framework designed to study treatment effect heterogeneity. Two directions warrant future studies. On the one hand, it is possible to extend our current framework to identify the top few subgroups instead of the best one. Take two subgroups for example; our goal can be formulated as finding the oracle treatment assignment that solves $\max_{\bm{e}}\ \mathbb{P}\big{(}\min\{\widehat{\tau}_{1},\widehat{\tau}_{2}\}\geq\max_{3\leq j\leq m}\widehat{\tau}_{j}\big{)}$, which can be achieved by revising the objective function as $\max_{\bm{e}}\min_{3\leq j\leq m}\max_{1\leq k\leq 2}\ G(\mathcal{S}_{k},\mathcal{S}_{j};e_{k},e_{j})$. On the other hand, the theoretical analysis presented in Section 6.3 is based on the assumption that the collected sample is independent and identically distributed. We hypothesize that this condition can be relaxed by utilizing refined concentration inequalities that incorporate martingales [13, 8]. Exploring these possibilities will be an avenue for future research.

### v3

@@ 1.1 Motivation
The existing literature in this research area mostly focuses on conducting retrospective analyses that employ observational or randomized experiment data. Even with large-scale observational data or carefully collected randomized experiment data, statistical bias in these analyses cannot be overlooked. On the one hand, in observational studies, statistical bias can arise due to potential violations of untestable causal assumptions [4, 5, 13, 18, 38, 47, 62, 71]. For example, one of the commonly imposed causal assumptions in practice is the unconfoundedness assumption, which states that conditional on measured confounders, the treatment assignment is as good as random [see 31, for example]. Due to the untestable nature of the unconfoundedness assumption and the possibility of unmeasured confounders in observational data, the validity of established causal conclusions under this assumption cannot be guaranteed [48, 62]. On the other hand, when analyzing classical randomized experiment data, although carrying out valid causal conclusion does not require imposing untestable causal assumptions, exploring treatment effect heterogeneity could still be susceptible to the winner’s curse bias if seemingly promising heterogeneous treatment effects are selected from the data in an ad hoc fashion [2, 33, 63, 89].
@@ 1.2 Our contribution
We tackle three major challenges in our theoretical investigation, which arise as the data are sequentially collected, and hence, they are not statistically independent. First, leveraging upon martingale methods, we develop a general framework and show that consistent estimation of features of the potential outcome distributions is possible for a large class of adaptive experimental designs (Lemma C.1 and its corollary in the Supplementary Materials), provided that the allocation probability and the enrichment proportion are bounded away from zero. Importantly, this result does not require independence and allows the data to be collected in an adaptive manner. More broadly, this technical lemma not only unifies adaptive enrichment designs and response adaptive randomization but also applies to other adaptive design settings with objective functions different from ours. We expect the general consistency result to be of independent interest. The second challenge in our theoretical investigation is that the optimized treatment assignment rule, both in finite samples and asymptotically, may not be unique, rendering standard M-estimation proof techniques inapplicable to our setting. Nevertheless, building on the general consistency result (Lemma C.1), we are able to show that with probability approaching one, any empirically optimized treatment assignment rule will be arbitrarily close to one of the oracle allocations (Lemma C.2 in the Supplementary Materials). Such “set consistency” result, seems new in the literature on adaptive experimental design. Finally, we establish the asymptotic normality of our treatment effect estimators. This result is established with a martingale central limit theorem [36], and hence, it takes into account the adaptive nature of the proposed experimental design and the induced dependence. In this process, a key step is to show that the cumulative empirical treatment probability (or the enrichment proportion) also converges to its oracle counterpart, which in turn helps to verify that a conditional variance will stabilize asymptotically. See Section C.7 in the Supplementary Materials for details and additional discussions.
@@ 1.3 Existing literature
Our framework falls into the realm of frequentist response adaptive designs. Such design strategies can be roughly divided into response-adaptive randomization (RAR) design and adaptive enrichment design in the existing literature. RAR design often refers to the design strategy in which the treatment assignment probabilities are adapted during the experiment based on the accrued evidence in the outcomes, with the goal of simultaneously achieving the experimental objectives and preserving statistical inference validity [42, 75, 73]. The classical RAR framework often revises the treatment assignment probabilities at infinitely many stages, a design strategy we refer to as fully adaptive settings. In fully adaptive settings, a popular class of response-adaptive randomization designs is the doubly adaptive biased coin (DBCD) design. The early DBCD design can be found in [24], which has its root in Efron’s biased coin design [22]. The asymptotic properties of DBCD designs are studied in various works [43, 45, 92, 116]. Our work shares some connections with [92] in that we both incorporate an optimization perspective into the problem of finding optimal treatment allocation, although our design objectives and the implementation of the optimal treatment allocation differ. Other than the fully adaptive settings, existing RAR designs also accommodate multi-stage settings [70, 93, 112]. Such multi-stage designs propose to revise the treatment assignment probability by minimizing the asymptotic variance of the average treatment effect estimator. Nevertheless, this design is carried out in two stages and is not designed to identify treatment effect heterogeneity.
Instead of relying solely on the outcome variable to optimize for the experimental goals in RAR, one may further incorporate covariate information. RAR that further incorporate covariate information is known as covariate-adjusted response-adaptive (CARA) designs [9, 78, 111, 93]. Early work in [106] proposes to balance covariates based on the biased coin design. [46] propose a family of CARA designs that could account for both efficiency and ethics. [115] generalizes CARA to incorporate semiparametric estimates. Some related CARA designs are also discussed in [60, 94, 113, 112]. We note that much of the existing work on response-adaptive randomization designs and covariate-adjusted response-adaptive aims to optimize the estimation efficiency of the overall treatment effect but is not tailored to study treatment effect heterogeneity [41, 76].
@@
As the data are sequentially collected using our design strategy, the toolkit we adopted to ensure the validity of statistical inference (martingale limit theories in particular) has also been used in the existing literature. For example, [61, 35, 107, 108] focus on analyzing adaptively collected data either from adaptive randomized experiments or online policy learning, different from our goal in designing an adaptive data collection mechanism.
Our work is connected to the literature on learning optimal policy. This area of literature is referred to by different names, such as choosing the best policy (e.g., [51]) and identifying the best arm (e.g., [84]). Other related work includes [7], [28], and [52]. The differences between our work and theirs arise from the goals and constraints. In our work, the allocation ratios between control and treatment in different subgroups are determined by solving an optimization problem with the goal of maximizing the correct selection probability under the constraints of maintaining a non-zero propensity score and having limited resources to conduct the experiment. In contrast, [51] adopt an allocation ratio between two arms that is asymptotically 1:1 directly enforced by an “exploration sampling” constraint. In [84], their top-two series algorithms by default have the allocation ratio between the “top two” arms set as 1:1 and instead focus on using posterior probability to update which two arms are “top two” on the fly. In our work, the optimal allocation ratio between two arms is not necessarily 1:1. Our work involves situations where implementing a treatment is costly, and there is an overall budget for how many treatments can be deployed. Such a budget constraint may impact the optimal allocation ratio. On the other hand, even with an unlimited budget, in our work, the optimal allocation ratio is not necessarily 1:1. Instead, it depends on the variances and relative mean gap between the two arms across different subgroups, which is rooted in our goal of maximizing the correct selection probability for the best subgroup. The differences in the goals and constraints between our work and the literature not only affect the treatment allocation and statistical methodology but also introduce additional methodological and technical challenges.
Furthermore, we would like to point out an additional delicate difference between our work and the broad learning and bandit literature. In most bandit problems, due to their prevalent applications within online recommendation, the most limited resource is a sample of users (experimentation unit), whereas the cost of various treatments is negligible. In our work, we consider situations where implementing the treatment can be costly per sample of a user (experimentation unit). For example, in clinical settings, randomized experiments can be expensive due to the high costs associated with treatment medication. Therefore, the resource constraints in our problem remain at the experimentation unit level and also at the level of how many treatments can be provided according to a budget.
Our work is also connected to the broad class of “contextual bandit” problems (e.g., [21] and subsequent work). Contextual bandit problems aim to learn an optimal policy that maps a given covariate to the best arm for that covariate, where the metric is either to minimize cumulative regret or to minimize simple regret. In particular, the metric of minimizing a simple regret is connected to our metric of maximizing correct selection probability, although our goal is different from that of the contextual bandit problems. Instead of the goal of contextual bandit problems to identify the best arm for each subgroup (each covariate value), our work aims to identify the subgroup that benefits the most from a single new treatment under resource constraints. The optimization involved in solving for the optimal resource allocation is, therefore, different and requires different techniques to solve and analyze.
@@ 4.1 Oracle treatment allocation in response-adaptive randomization designs
This equation suggests that the required number of participants in the treatment arm of subgroup $j$ is reduced when it is relatively easier to distinguish subgroup $j$ from the best subgroup. This occurs when there is a larger difference between $\tau_{j}$ and $\tau_{1}$ or when the variance of subgroup $j$ is higher. To provide a clearer understanding, we consider a simple scenario with $m=3$. In Figure 1, we plot the relationship between $e^{*}_{2}$, $\tau_{2}$, and $\sigma_{2}(1)^{2}$.
@@ 5.1 Oracle subgroup enrichment proportions
In enrichment designs, the treatment allocation probabilities are often fixed; we thus denote the rate function as a function of subgroup enrollment proportions:
@@ 5.2 Proposed adaptive enrichment designs
Applying the same reasoning as discussed in Section 4.3, we calculate the estimated optimal subgroup enrichment proportion by subtracting the proportion of enrolled subjects in the previous stages. If Eq (11) suggests multiple possible designs, one can choose an enrichment proportion that is closest to the uniform allocation of $1/m$ or a higher enrichment proportion for the subgroup with a larger treatment effect.
@@
Construct the final subgroup treatment effect estimator $\widehat{\tau}_{j}^{AE}$ and the variance estimator $\widehat{\mathbb{V}}_{j}^{AE}$ under the adaptive enrichment (“AE”) design. These two estimators take the same form as in Eq (5) and Eq (6).
@@
$\Bigg{[}\widehat{\tau}^{\texttt{AE}}_{j^{*}}\pm\Phi^{-1}(1-\alpha/2)\cdot\sqrt{\widehat{\mathbb{V}}^{\texttt{AE}}_{j^{*}}/N}\Bigg{]}.$
@@ 6.3 Comparison with completely randomized experiments
$\begin{cases}\mathbb{V}_{j}(e_{j}^{*})\leq\mathbb{V}_{j}(c_{1}),\quad&\text{if}\ (\tau_{j}-\tau_{1})^{2}-\frac{1}{e_{1}^{*}(1-e_{1}^{*})}\leq\frac{1}{c_{1}^{2}},\\ \mathbb{V}_{j}(e_{j}^{*})>\mathbb{V}_{j}(c_{1}),\quad&\text{if}\ (\tau_{j}-\tau_{1})^{2}-\frac{1}{e_{1}^{*}(1-e_{1}^{*})}>\frac{1}{c_{1}^{2}}.\end{cases}$
@@ 7.2 Synthetic data generation and simulation setup
Under each design setup, we compare our proposed design strategy with other conventional designs as summarized in Table 2. In Table 2 , the “complete randomization” design refers to setting the treatment assignment probability $e_{tj}=\frac{1}{2}$ in all experimental stages, for $t=1,\ldots,T,\ j=1,\ldots,m$. The “Neyman allocation” refers to setting the treatment assignment probability $e_{tj}=\frac{\sigma_{1j}}{\sigma_{1j}+\sigma_{0j}}$. The “Proposed design combined with DBCD” refers to enhancing our proposed design with the doubly adaptive biased coin design (DBCD) in [109]. The DBCD design is a response-adaptive randomization design that targets the current treatment allocation towards the optimal treatment allocation. As we consider assigning treatments to multiple subgroups, we use the DBCD design to target the optimal treatment allocation in each subgroup separately. Our implementation is summarized as follows:
  1. (1)
    At Stage $t$, obtain optimal treatment allocations $\widehat{e}_{t,j}^{*}$, $j=1,\ldots,m$ by solving the optimization problem as in Section 4.2. Calculate current treatment allocation up to Stage $t-1$, denoted as $\widehat{e}_{t-1,j}$, $j=1,\ldots,m$.
  2. (2)
    For each subgroup $j$, calculate treatment allocation under the DBCD design proposed in [43]:
    $\psi_{t,j}(\widehat{e}_{t,j}^{*},\ \widehat{e}_{t-1,j})=\frac{\widehat{e}_{t,j}^{*}(\frac{\widehat{e}_{t,j}^{*}}{\widehat{e}_{t-1,j}})^{\gamma}}{\widehat{e}_{t,j}^{*}(\frac{\widehat{e}_{t,j}^{*}}{\widehat{e}_{t-1,j}})^{\gamma}+(1-\widehat{e}_{t,j}^{*})(\frac{1-\widehat{e}_{t,j}^{*}}{1-\widehat{e}_{t-1,j}})^{\gamma}},$
    where $\gamma\in[0,\infty)$ is a tuning parameter.
  3. (3)
    At Stage $t$, we assign treatments with probability $\psi_{t,j}(\widehat{e}_{t,j}^{*},\ \widehat{e}_{t-1,j})$ in each subgroup $j$.
The “equal enrichment design” refers to the design that sets the enrichment proportion as $p_{tj}=\frac{1}{m}$ across all the experimental stages. The “adaptive enrichment with combination testing” approach is a method that distributes the type I error rate across experimental stages. Based on the computed type I error rate each stage aims to reach, the corresponding enrichment proportions can be estimated. We implement the combination testing approach using R package rpact [57].
@@
(c) Neyman allocation
(d) Proposed design combined with DBCD
@@ 7.3 Synthetic case study results
In Figures 4 to 7, we compare our proposed response-adaptive randomization design with the other conventional designs in the fully adaptive setting and the multi-stage setting. We summarize our simulation results from three aspects, following the order outlined at the start of Section 7
@@
Lastly, in the RAR setting, combining our proposed design with the DBCD design can further enhance the finite sample performance. Figures 4 and 5 demonstrate that DBCD design can enhance the performance of our method in finite samples. When using the DBCD design to target our actual treatment allocation towards the optimal treatment allocation, the design strategy exhibits a smaller estimation bias for the best subgroup treatment effect and an increase in the correct selection probability. As the sample size increases, the performances of our proposed design and the DBCD-enhanced design tend to converge. In Figure 6, we provide an additional simulation study to highlight the broad benefits of DBCD design in enhancing the finite sample performance of various designs. We compare three designs: (i) complete randomization + DBCD, (ii) Neyman allocation + DBCD, and (iii) Proposed RAR design + DBCD. We use “+DBCD” to indicate that the DBCD is applied to each design to guide the treatment allocation closer to the optimal treatment allocation. Figure 6 shows that DBCD design generally improves the finite-sample performance of these design strategies by effectively targeting the current treatment allocation towards the optimal treatment allocation.
Figure 4: The estimated treatment effects and the associated standard errors in the four clothing categories under the complete randomization design, Neyman allocation, our proposed response-adaptive randomization design, and our proposed design in combination with the DBCD design in the fully adaptive setting ($N=400$).
Figure 5: Comparison of the proposed response-adaptive randomization design, the complete randomization design, Neyman allocation, and our proposed design in combination with the doubly adaptive biased coin design under the fully adaptive setting. (A) shows the sample size comparison under various correct selection probability levels. (B) shows the $\sqrt{N}$-scaled winner’s curse bias comparison with respect to different sample sizes.
Figure 6: Comparison of the proposed response-adaptive randomization design, the complete randomization design, and the Neyman allocation in combination with the doubly adaptive biased coin design under the fully adaptive setting. (A) shows the sample size comparison under various correct selection probability levels. (B) shows the $\sqrt{N}$-scaled winner’s curse bias comparison with respect to different sample sizes.
@@ 8 Discussion
In this manuscript, we propose a unified adaptive experimental framework designed to study treatment effect heterogeneity. Three directions warrant future studies. First, it is possible to extend our current framework to identify the top few subgroups instead of the best one. Take two subgroups for example; our goal can be formulated as finding the oracle treatment assignment that solves $\max_{\bm{e}}\ \mathbb{P}\big{(}\min\{\widehat{\tau}_{1},\widehat{\tau}_{2}\}\geq\max_{3\leq j\leq m}\widehat{\tau}_{j}\big{)}$, which can be achieved by revising the objective function as $\max_{\bm{e}}\min_{3\leq j\leq m}\max_{1\leq k\leq 2}\ G(\mathcal{S}_{k},\mathcal{S}_{j};e_{k},e_{j})$. Second, the theoretical analysis presented in Section 6.3 is based on the assumption that the collected sample is independent and identically distributed. We hypothesize that this condition can be relaxed by utilizing refined concentration inequalities that incorporate martingales [15, 10]. Exploring these possibilities will be an avenue for future research. Third, our design considers the setting when outcomes are observed without delay. The significance of incorporating delayed responses in adaptive trials has been discussed and recognized in various adaptive design literature, including [79] and [73]. Some existing adaptive experiment methods and theoretical results related delayed outcomes have been discussed under the urn models [101, 100, 8, 43, 110], the doubly-adaptive biased coin design (DBCD) framework [109, 44, 54], and in the group-sequential settings [37, 86, 30]. We hope to extend our proposed design and adjust it for delayed outcomes in our future research.